OpenStudy (anonymous):

8x^2y^3 -------- please help! 6yx^0

OpenStudy (anonymous):

\[\frac{4x^2y^2}{3}\]

OpenStudy (anonymous):

i dont get it please explain

OpenStudy (anonymous):

Ok first thing to do is to simplify the fraction \[\frac{8}{6}\] which is \[\frac{4}{3}\] so now you have \[\frac{4x^2y^3}{3x^0y}\]

OpenStudy (anonymous):

Now when you are dividing numbers with exponents you subtract the top exponent from the bottom exponent. So now you have \[\frac{4x^{2-0}y^{3-1}}{3}\]

OpenStudy (anonymous):

Which leaves you with \[\frac{4x^2y^2}{3}\]

OpenStudy (anonymous):

oh i see thanks!